# PAC-Bayes Analysis for Recalibration in Classification

### v1

@@ 7 Conclusion and limitations
This paper provides the first comprehensive analysis of the generalization error and estimation bias of the ECE in both binary and multiclass classifications, resulting in several non-asymptotic bounds for them and the practical optimal bin size. Moreover, we develop a new generalization-aware recalibration algorithm based on our PAC-Bayes bound. Our analysis also reveals the slow convergence of the ECE through binning, which is a fundamental limitation. As discussed in Section 3.1, the binning method cannot leverage the underlying smoothness of the data distribution. Our numerical results suggest that the limited number of test data points makes the numerical evaluation of the ECE unstable in practice. We believe that other nonparametric methods used to estimate the CE or TCE may resolve this issue. An investigation of this point is our important future work.
Acknowledgements
MF was supported by RIKEN Special Postdoctoral Researcher Program. MF was supported by JST, ACT-X Grant Number JPMJAX210K, Japan. FF was supported by JSPS KAKENHI Grant Number JP23K16948. FF was supported by JST, PRESTO Grant Number JPMJPR22C8, Japan.
Appendix A Proofs and discussion of Section 3
@@ A.1 Additional preliminary
[2] We say that a function $f:\mathcal{X}\to\mathbb{R}$ has the bounded difference property if for some nonnegative constants $c_{1},\dots,c_{n}$,
@@
A.2 Proof of Theorem 2
A.2.1 Reformulations of the ECE to loss functions
Here, we introduce the reformulations of the ECE, which is the essential technique in our proof.
@@ A.2.1 Reformulations of the ECE to loss functions
$\mathrm{ECE}(f_{w},S_{\mathrm{tr}})\coloneqq\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}|,$ (19)
which follows immediately from the definitions. Here $\widehat{S}_{\mathrm{tr}}$ is the empirical distribution of the training dataset.
Recall the conditional expectation of $f_{w}$ given bins
$f_{\mathcal{I}}(x)\coloneqq=\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)\in I_{i}}.$ (20)
@@
$=\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})\mathbb{E}|\mathbb{E}[Y|f_{w}(X)\in I_{i}]-\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]|,$ (22)
@@
$\mathrm{ECE}(f_{w},S_{e})=\sum_{i=1}^{B}\frac{|I_{i}|}{n_{e}}|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}|=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim\widehat{S}_{\mathrm{tr}}}(\mathbbm{1}_{Y=C}-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}},$ (24)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}f_{w}(x_{m})$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}\mathbbm{1}_{y_{m}=C}$. As for the conditional function, we define
$f_{\mathcal{I}}^{C}(x)\coloneqq\sum_{i=1}^{B}f_{I_{i}}^{C}(x)\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}=\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)_{C}|f_{w}(X)_{C}\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}.$ (25)
@@
A.2.2 Preliminary of the proof of Theorem 2
Next, we introduce the assumption, which is used in existing analysis Gupta and Ramdas [11] and not necessarily in our proof;
@@ A.2.2 Preliminary of the proof of Theorem 2
Given $W=w$, $f_{w}(x)$ is absolutely continuous with respect to the Lebesgue measure.
This assumption means that $f_{w}(x)$ has a probability density, and $f_{w}(x)$, it is satisfied without loss of generality as elaborated in Appendix C in Gupta and Ramdas [11]. Although this assumption is not required for our main results, if we assume this, we can improve the coefficient of the upper bound. Thus in the below, we provide the proof with and without this assumption simultaneously.
Thanks to this assumption, $f_{w}(X)$ takes distinct values almost surely, so for example, the situation when all training samples take the same predicted probability will be circumvented. When considering the multiclass setting with $CE_{K}$, this property plays an important role in eliminating the curse of dimensionality, see Appendix A.6.
We start from the slightly different version of Theorem 2. Then we extend it  2.
@@
$\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}.$ (27)
@@
$\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{\lambda^{2}}{2n}}{\lambda}.$ (28)
@@
We first prove the binary setting. To derive the PAC Bayes bound, we transform the loss using Eq. (19) and (21), we have
@@
where we used the triangle inequality $||a|-|b||\leq|a-b|$ for the first inequality and set $l_{i}(z)=(y-f_{W}(x))\cdot\mathbbm{1}_{f_{W}(x)\in I_{i}}$.
@@
Combined with the fact that $|y^{\prime}_{m^{\prime}}-f_{w}(x^{\prime}_{m^{\prime}})|\leq 1$, we obtain Eq. (34). Note that by Assumption 3, $\{f_{w}(x_{m})\}_{m=1}^{n}$ in $x_{m}\in S^{\prime}$ takes the distinct values almost surely and in the above discussion, we do not consider the case when $b/B=f_{w}(x_{m})$ for some $b$ holds, which means that the predicted probability is just the value of the boundary of bins.
@@
$\mathbb{E}_{S^{\prime}}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\prod_{m=1}^{n}\mathbb{E}_{Z^{\prime}_{m}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime\prime}}l_{i}(Z^{\prime\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m}^{\prime})\right]}$
@@
$\mathbb{E}_{S^{\prime}}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\prod_{m=1}^{n}\mathbb{E}_{Z^{\prime}_{m}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime\prime}}l_{i}(Z^{\prime\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m}^{\prime})\right]}\leq 2^{B}e^{\frac{2t^{2}}{n}}.$ (37)
@@
A.2.3 Proof of Theorem 2 (Generalization of the ECE under the training dataset)
Now we present the proof of Theorem 2

### v2

@@ 7 Conclusion and Limitations
In this paper, we analyzed the generalization error and estimation bias of the ECE in multiclass classification for the first time, resulting in several non-asymptotic bounds and the practical optimal bin size. Moreover, we developed a new generalization-aware recalibration algorithm based on our PAC-Bayes bound. Our analysis also revealed the slow convergence of the ECE through binning, which is a fundamental limitation. As discussed in Section 3.2, the binning method cannot leverage the underlying smoothness of the data distribution. Our numerical results suggest that the limited number of test data points makes the numerical evaluation of the ECE unstable in practice. We believe that other nonparametric methods for estimating the CE or TCE may address the limitations identified in this study. Investigating the effectiveness of such methods constitutes an important direction for future work. Another key avenue for future research is to extend the theoretical analysis to cases where alternative methods for estimating the TCE, such as UMB, are employed.
Acknowledgments
MF was previously supported by the RIKEN Special Postdoctoral Researcher Program and JST ACT-X Grant Number JPMJAX210K, Japan. MF is currently supported by KAKENHI Grant Number 25K21286. FF was supported by JSPS KAKENHI Grant Number JP23K16948. FF was supported by JST, PRESTO Grant Number JPMJPR22C8, Japan.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Remark about the binary setting
Here we introduce the calibration for binary classification. Although, the primary goal of our study is the multiclass setting, our theory can handle for both the binary and multicalss settings.
Let $\mathcal{Y}=\{0,1\}$ and let $f_{w}:\mathcal{X}\to[0,1]$ be a model, where the output corresponds to the model’s confidence that the label is $1$. In the context of calibration, it is expected that not only $f_{w}$ has high accuracy, but also its predictive probability aligns well in the actual label frequency $P(Y|f_{w}(X))$. A model is well-calibrated [Widmann et al., 2021, Gupta et al., 2020] when $P(Y|f_{w}(X))=f_{w}(X)$ holds; however, it is difficult to confirm this. Therefore, an alternative metric called as the calibration error (CE) [Widmann et al., 2021, Gupta et al., 2020] is typically used. Given $W=w$, the CE is defined as
Unfortunately, evaluating the CE is still infeasible because $\mathbb{E}[Y|f_{w}(X)]$ is intractable. To resolve this issue, we evaluate the calibration performance by constructing the estimator of the CE. The widely used estimator is the expected calibration error (ECE), where binning is used to estimate $\mathbb{E}[Y|f_{w}(X)]$ [Guo et al., 2017, Zadrozny and Elkan, 2001, 2002]. The ECE is calculated by partitioning the range of the predictive probability $[0,1]$ into $B$ intervals =$\mathcal{I}=\{I_{i}\}_{i=1}^{B}$ (called bins) and averaging within each bin using an evaluation dataset $S_{e}\coloneqq\{z_{m}\}_{m=1}^{n_{e}}\in\mathcal{Z}^{n_{e}}$, where we assume $n_{e}\geq 2B$. That is, the ECE is defined as
Similarly to the multiclass setting, we define the total bias of CE as $\mathrm{Bias}(\eta_{v}(f_{w}),S_{e},\mathrm{CE})\coloneqq|\mathrm{CE}(\eta_{v}(f_{w}))-\mathrm{ECE}(\eta_{v}(f_{w}),S_{e})|$ for binary classification.
Appendix B Proofs and discussion of Section 3
@@ B.1 Additional preliminary
[Boucheron et al., 2013] We say that a function $f:\mathcal{X}\to\mathbb{R}$ has the bounded difference property if for some nonnegative constants $c_{1},\dots,c_{n}$,
@@
The following theorem is the most basic form of the PAC-Bayes generalization bounds [Alquier et al., 2016].
Theorem 4.
B.2 Reformulations of the ECE to loss functions
Here, we introduce the reformulations of the ECE, which is the essential technique in our proof. Note that when focusing on the TCE of the multiclass setting, its properties are almost the same as those of the binary classification. This is because TCE only considers the top label and by regarding the top label corresponding to label $1$ in the binary classification and other labels as 00 in the binary classification, then it is clear that the ECE of the multiclass and binary setting are almost identical.
@@ B.2 Reformulations of the ECE to loss functions
$\mathrm{ECE}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim S_{e}}(Y-f_{w}(X))\cdot\mathbbm{1}_{f_{w}(X)\in I_{i}}|,$ (32)
which follows immediately from the definitions. Here $\mathbb{E}_{S_{e}}$ is the empirical expectation by the evaluation dataset.
Recall that the conditional expectation of $f_{w}$ given bins is defined as
@@
$=\sum_{i=1}^{B}P(f_{w}(X)\in I_{i})|\mathbb{E}|\mathbb{E}[Y|f_{w}(X)\in I_{i}]-\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]|,$ (35)
@@
$\mathrm{ECE}(f_{w},S_{e})=\sum_{i=1}^{B}\frac{|I_{i}|}{n_{e}}|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}|=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim S_{e}}(\mathbbm{1}_{Y=C}-f_{w}(X)_{C})\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}|,$ (37)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}f_{w}(x_{m})_{C}$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}\mathbbm{1}_{y_{m}=C}$. As for the conditional function, we define
@@
B.3 Proof of Theorem 2 (Bias under the training dataset)
Here we present a generalized version of Theorem 2 that includes the binary classification. After that, we show the proof of Theorem 1, which can be directly obtained from Theorem 2.
To simplify the notation, we express $n_{\mathrm{tr}}=n$.
Theorem 5.
Under Assumption 1, for any fixed prior $\pi$, which is independent of $S_{\mathrm{tr}}$, and posterior distribution $\rho$ over $W$ and any positive constant $\varepsilon\in[0,1]$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{tr}}$, we have
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}\quad(\text{Binary classification}),$ (40)
$\mathbb{E}_{\rho}\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}\quad(\text{Multiclass classification}).$ (41)
Proof.
Binary setting
We start from the binary classification setting. To analyze the bias caused by binning, following Futami and Fujisawa [2024], we define the conditional expectation of $f_{w}$ given bins as $f_{\mathcal{I}}(x)\coloneqq\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)\in I_{i}}$. With this definition, we decompose the estimation bias as follows:
$\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})$ $=|\mathrm{CE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{te}})|\leq|\mathrm{CE}(f_{w})-\mathrm{CE}(f_{\mathcal{I}})|+|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{w},S_{\mathrm{te}})|.$ (42)
where the first term is called as the binning bias, which arises from nonparametric estimation via binning, and the latter as the statistical bias caused by estimation on finite data points. We then evaluate these terms separately.
As for the first term, we show that
holds. Note that $P(Y=1|f_{w}(x))=\mathbb{E}[Y|f_{w}(x)]$ holds for the binary classification. First, we can show that
$\mathrm{CE}(f_{\mathcal{I}})\leq\mathrm{CE}(f_{w}),$ (44)
$\quad+\mathbb{E}[f_{w}(X)|f_{w}(X)\in I_{i}]-f_{w}(X)||f_{w}(X)\in I_{i}]$
Then we can show that
where we used the Lipschitz continuity of the function and the fact that we split the function with equal width $1/B$. Now, we have
We will bound the second term. For this purpose, we use Theorem 6 in Appendix B.4, this concludes the proof for the binary setting. ∎
Proof.
Multiclass setting
Next, for the multiclass setting, similarly to the binary case, Similarly to the binary setting, we define the conditional expectation of $f_{w}$ given bins as $f_{\mathcal{I}}^{C}(x)\coloneqq\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)_{C}|f_{w}(X)_{C}\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}$. With this definition, we decompose the estimation bias as follows:
$\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{TCE})=|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{te}})|\leq|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f_{\mathcal{I}}^{C})|+|\mathrm{TCE}(f_{\mathcal{I}}^{C})-\mathrm{ECE}(f_{w},S_{\mathrm{te}})|.$ (48)
As for the multiclass setting, similarly to the binary case, we have
$|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f^{C}_{\mathcal{I}})|\leq\mathbb{E}|P(Y=C|f_{w}(X)_{C})-P(Y=C|f_{\mathcal{I}}^{C}(X)|+\mathbb{E}|f_{w}(X)_{C}-f_{\mathcal{I}}^{C}(X)|$ (50)
by the almost identical proof. We then upper-bound them
$|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f^{C}_{\mathcal{I}})|\leq\mathbb{E}|P(Y=C|f_{w}(X)_{C})-P(Y=C|f_{\mathcal{I}}^{C}(X)|+\mathbb{E}|f_{w}(X)_{C}-f_{\mathcal{I}}^{C}(X)|\leq\frac{L}{B}+\frac{1}{B}$ (51)
$\mathbb{E}_{\rho}|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\!$
Then using Theorem 6 in Appendix B.4 for the second term, this concludes the proof. ∎
B.4 Auxiliary result about the statistical bias
Here we present the auxiliary result, which is used in Appendix B.3.
First, we introduce the assumption, which is used in existing analysis Gupta and Ramdas [2021] and not necessarily in our proof;
@@ B.4 Auxiliary result about the statistical bias
Given $W=w$, $f_{w}(x)_{C}$ is absolutely continuous with respect to the Lebesgue measure.
As for the binary setting, we assume that $f_{w}(x)$ is absolutely continuous with respect to the Lebesgue measure. This assumption means that $f_{w}(x)$ has a probability density, it is satisfied without loss of generality as elaborated in Appendix C in Gupta and Ramdas [2021]. Although this assumption is not required for our main results, if we assume this, we can improve the coefficient of the upper bound. Thus in the below, we provide the proof with and without this assumption simultaneously.
Thanks to this assumption, $f_{w}(X)$ takes distinct values almost surely, so for example, the situation when all training samples take the same predicted probability will be circumvented. When considering the multiclass setting with $\mathrm{CE}_{K}$, this property plays an important role in eliminating the curse of dimensionality, see Appendix B.9.
@@
$\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}\quad(\text{Binary classification}),$ (53)
$\mathbb{E}_{\rho}|\mathrm{TCE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n}}{\lambda}\quad(\text{Multiclass classification}).$ (54)
@@
$\mathbb{E}_{\rho}|\mathrm{CE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{\lambda^{2}}{2n}}{\lambda}\quad(\text{Binary classification}),$ (55)
$\mathbb{E}_{\rho}|\mathrm{TCE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S_{\mathrm{tr}})|\leq\frac{\mathrm{KL}(\rho\|\pi)+B\log 2+\log\frac{1}{\varepsilon}+\frac{\lambda^{2}}{2n}}{\lambda}\quad(\text{Multiclass classification}).$ (56)
@@
Binary setting
To derive the PAC Bayes bound, we transform the loss using Eq. (32) and (34), we have
@@
where we used the triangle inequality $||a|-|b||\leq|a-b|$ for the first inequality and set $l_{i}(z)=(y-f_{W}(x))\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}$.
@@
Combined with the fact that $|y^{\prime}_{m^{\prime}}-f_{w}(x^{\prime}_{m^{\prime}})|\leq 1$, we obtain Eq. (62). Note that by Assumption 3, $\{f_{w}(x_{m})\}_{m=1}^{n}$ in $x_{m}\in S$ takes the distinct values almost surely and in the above discussion, we do not consider the case when $b/B=f_{w}(x_{m})$ for some $b$ holds, which means that the predicted probability is just the value of the boundary of bins.
@@
$\mathbb{E}_{S}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\prod_{m=1}^{n}\mathbb{E}_{Z^{\prime}_{m}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime\prime}}l_{i}(Z^{\prime\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m}^{\prime})\right]}$
@@
$\mathbb{E}_{S}e^{t\sum_{i=1}^{B}|g(i,Z^{\prime}_{m})|}$ $\leq\sum_{v_{1},\dots,v_{B}=0,1}\prod_{m=1}^{n}\mathbb{E}_{Z^{\prime}_{m}}e^{t\sum_{i=1}^{B}(-1)^{v_{i}}\left[\mathbb{E}_{Z^{\prime\prime}}l_{i}(Z^{\prime\prime})-\frac{1}{n}\sum_{m=1}^{n}l_{i}(Z_{m}^{\prime})\right]}\leq 2^{B}e^{\frac{2t^{2}}{n}}.$ (65)
@@
Multiclass setting
As for the ECE of the multiclass setting, from Eq. (37) and Eq. (39), we can proceed the proof exactly the same way as the binary setting. By using the triangular inequality similarly to Eq. (57) and setting $l_{i}(Z)=(\mathbbm{1}_{Y=C}-f_{w}(X)_{C})\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}$, we have
$|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\leq\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim\mathcal{D}}l_{i}(Z)-\mathbb{E}_{Z\sim S_{\mathrm{tr}}}l_{i}(Z)\right|.$ (70)
By using McDiarmid’s inequality to evaluate the exponential moment, we obtain
$\mathbb{E}_{S,\pi}e^{t|\mathrm{TCE}(f_{\mathcal{I}})-\mathrm{ECE}(f_{W},S)|}\leq\mathbb{E}_{S,\pi}e^{\lambda\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim\mathcal{D}}l_{i}(Z)-\mathbb{E}_{Z\sim\hat{S}_{\mathrm{tr}}}l_{i}(Z)\right|}\leq 2^{B}e^{\frac{2\lambda^{2}}{n}}$ (71)
Then by using the Fubini theorem and Donsker-Valadhan inequality as in the case of the binary classification, we get the bound. ∎
B.5 Proof of Theorem 1 (Total bias analysis)
Now we present the proof of Theorem 1. The proof is almost identical to that of Theorem 2 shown in Appendix B.3.
To analyze the bias caused by binning, following Futami and Fujisawa [2024], we define the conditional expectation of $f_{w}$ given bins as $f_{\mathcal{I}}^{C}(x)\coloneqq\sum_{i=1}^{B}\mathbb{E}[f_{w}(X)_{C}|f_{w}(X)_{C}\in I_{i}]\cdot\mathbbm{1}_{f_{w}(x)_{C}\in I_{i}}$. With this definition, we decompose the estimation bias as follows:
$|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{te}})|\leq|\mathrm{TCE}(f_{w})-\mathrm{TCE}(f_{\mathcal{I}}^{C})|+|\mathrm{TCE}(f_{\mathcal{I}}^{C})-\mathrm{ECE}(f_{w},S_{\mathrm{te}})|.$ (72)
where the first term is called as the binning bias, which arises from nonparametric estimation via binning, and the latter as the statistical bias caused by estimation on finite data points. We then evaluate these terms separately.
As for the binning bias, using Assumption 1, it can be bounded by $(1+L)/B$ because prepared bins divide the interval $[0,1]$ into equal widths.
Next, we evaluate the statistical bias. We first derive the following reformulation:
$\mathrm{ECE}(f_{w},S_{\mathrm{te}})=\sum_{i=1}^{B}|\mathbb{E}_{(X,Y)\sim S_{\mathrm{te}}}(\mathbbm{1}_{Y=C}-f_{w}(X)_{C})\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}|,$
where $\mathbb{E}_{S_{\mathrm{te}}}$ is the empirical expectation for $S_{\mathrm{te}}$. This reformulation is derived in Eq. (37). As for the TCE, we get a similar relation from Eq. (39).
By using the triangular inequality similarly to Eq. (57) and setting $l_{i}(Z)=(\mathbbm{1}_{Y=C}-f_{w}(X)_{C})\cdot\mathbbm{1}_{f_{w}(X)_{C}\in I_{i}}$, we have
$|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|\leq\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim\mathcal{D}}l_{i}(Z)-\mathbb{E}_{Z\sim S_{\mathrm{te}}}l_{i}(Z)\right|.$ (73)
$\mathbb{E}_{S_{\mathrm{te}}}e^{\lambda\sum_{i=1}^{B}\left|\mathbb{E}_{Z\sim\mathcal{D}}l_{i}(Z)-\mathbb{E}_{Z\sim S_{\mathrm{te}}}l_{i}(Z)\right|}\leq 2^{B}e^{\frac{2\lambda^{2}}{n_{\mathrm{te}}}}$ (74)
for any $\lambda\in\mathbb{R}$ (see Eq. (66)). Combining this result with the Markov inequality, we can upper bound the statistical bias.
B.6 Generalization of the ECE under the training dataset
Here we show the generalization error bound for the ECE, which is useful when analyzing the recalibration algorithm.